OpenStudy (anonymous):

A die with 60 sides is tossed once. Find the probability that a number divisible by 30 is rolled. Round answer to 2 decimal places.

OpenStudy (anonymous):

number divisible by 30: 0, 30, 60 3/ 60 = 1/20 = 0.05